# Athena: Enhancing Multimodal Reasoning with Data-efficient Process Reward Models

### v2

@@ Appendix B Related Work
Test-time scaling.  While scaling data and model parameters during training has been extensively studied (Kaplan et al., 2020; Hoffmann et al., 2022), scaling computation at test-time has garnered interest more recently (Snell et al., 2025; Setlur et al., 2025). TTS enables language models to leverage additional computational resources when faced with challenging questions. Our study primarily focuses on parallel scaling (Zhang et al., 2025a), wherein language models generate multiple outputs simultaneously and aggregate them into a final answer. This aggregation can involve selecting the most common outcome (Wang et al., 2023) or using reward models (Snell et al., 2025) to rank solutions, selecting the one with the highest reward, as demonstrated in our Best-of-N settings.
Process reward models.  Reward models are crucial in reinforcement learning (Guo et al., 2025; Shao et al., 2024; Wang et al., 2024b; Ouyang et al., 2022; Bai et al., 2022; Dong et al., 2024) and test-time scaling (Setlur et al., 2025; Lightman et al., 2024; Wang et al., 2024b; Cobbe et al., 2021; Snell et al., 2025). Outcome reward models (Cobbe et al., 2021) assign a scalar value to question-response pairs, whereas process reward models evaluate each step, typically achieving superior performance and generalization (Lightman et al., 2024; Luo et al., 2024; Setlur et al., 2025; Li & Li, 2025). While better performance and generalization are achieved, collecting data to train PRMs is challenging. PRM800K (Lightman et al., 2024) is the first open-source process supervision dataset annotated by humans, but scaling it is challenging due to the time and skill required for annotating reasoning steps. Math-Shepherd (Wang et al., 2024b) proposes an automated method for estimating intermediate step correctness using Monte Carlo estimation, though it generates some incorrect labels and demands extensive computational resources. Our work explores strategies to reduce computational costs while improving process label accuracy.
@@ C.1 Choices of Completers
Table 6: The accuracy of estimated process labels with Different completers combination. In Two Completers, we use consistency filter introduced in Sec. 2.2. $\phi_{w}^{1}$, $\phi_{w}^{2}$, $\phi_{s}^{1}$, and $\phi_{w}^{2}$ denote Mistral-7B-Instruct, Qwen2-7B-Instruct, Qwen2.5-72B-Instruct and Llama-3-70B-Instruct, respectively.
@@ C.2 Benchmarks
MathVista (Lu et al., 2024b) Testmini 1000 Accuracy
MathVision (Wang et al., 2024a) Full 3040 Accuracy
MathVerse (Zhang et al., 2024) Vision-Only 788 Accuracy
DynaMath (Zou et al., 2025) Full 5050 Accuracy
WeMath (Qiao et al., 2025) Testmini 1740 Score (Strict)
LogicVista (Xiao et al., 2024) Full 448 Accuracy
MMMU (Yue et al., 2024) Validation 900 Accuracy
GSM8K (Cobbe et al., 2021) Test 1319 Accuracy
MATH Hendrycks et al. (2021) Test 5000 Accuracy
VisualProcessBench (Wang et al., 2025) - 2866 F1-score
@@ C.3 Reward Choice
In this paper, we consider three reward choices of Athena-PRM for aggregating step scores into a final score (Zhang et al., 2025b), including PRM-minium, PRM-last, and PRM-product. For all reasoning steps $\{a_{i}\}_{i=1}^{K}$, its corresponding reward for each step is $\{r_{i}\}_{i=1}^{K}$, the reward $r$ for each solution with different scoring methods are calculated as follows: 1) PRM-minium uses the minium reward across all steps, i.e., $r=\min\{r_{i}\}_{i=1}^{K}$; 2) PRM-last uses the reward of last step, i.e., $r=r_{K}$; 3) PRM-product uses the reward product of all steps, i.e. $r=\prod_{i=1}^{K}r_{i}$.
@@
Table 8: Best-of-8 performance comparison on the MathVista (Lu et al., 2024b) with Qwen2.5-VL-7B (Bai et al., 2025) with different reward choices: last step reward, minium reward across all steps and product of all rewards. Default settings are marked in gray.
@@ Appendix D Case Study
Figure 3: A case study from VisualProcessBench (Wang et al., 2025).
@@ 3. Angle properties:
Figure 4: A case study from VisualProcessBench (Wang et al., 2025).
@@
Figure 5: A case study from MathVista (Lu et al., 2024b) with Qwen2.5-VL-72B (Bai et al., 2025).
@@
Figure 6: A case study from MathVista (Lu et al., 2024b) with Qwen2.5-VL-72B (Bai et al., 2025).

### v3

@@ Appendix B Related Work
Test-time scaling.  While scaling data and model parameters during training has been extensively studied (kaplan2020scaling; hoffmann2022training), scaling computation at test-time has garnered interest more recently (snell2025scaling; setlur2025rewarding). TTS enables language models to leverage additional computational resources when faced with challenging questions. Our study primarily focuses on parallel scaling (tts-survey), wherein language models generate multiple outputs simultaneously and aggregate them into a final answer. This aggregation can involve selecting the most common outcome (wang2023selfconsistency) or using reward models (snell2025scaling) to rank solutions, selecting the one with the highest reward, as demonstrated in our Best-of-N settings.
Process reward models.  Reward models are crucial in reinforcement learning (deepseekr1; shao2024deepseekmath; wang2024mathshepherd; instruct_gpt; bai2022training; rlhf_workflow) and test-time scaling (setlur2025rewarding; lightman2024lets; wang2024mathshepherd; cobbe2021training; snell2025scaling). Outcome reward models (cobbe2021training) assign a scalar value to question-response pairs, whereas process reward models evaluate each step, typically achieving superior performance and generalization (lightman2024lets; OmegaPRM; setlur2025rewarding; li2025process_q_ranking). While better performance and generalization are achieved, collecting data to train PRMs is challenging. PRM800K (lightman2024lets) is the first open-source process supervision dataset annotated by humans, but scaling it is challenging due to the time and skill required for annotating reasoning steps. Math-Shepherd (wang2024mathshepherd) proposes an automated method for estimating intermediate step correctness using Monte Carlo estimation, though it generates some incorrect labels and demands extensive computational resources. Our work explores strategies to reduce computational costs while improving process label accuracy.
@@ C.1 Choices of Completers
Table 6: The accuracy of estimated process labels with Different completers combination. In Two Completers, we use consistency filter introduced in Sec. 2.2. $\phi_{w}^{1}$, $\phi_{w}^{2}$, $\phi_{s}^{1}$, and $\phi_{s}^{2}$ denote Mistral-7B-Instruct, Qwen2-7B-Instruct, Qwen2.5-72B-Instruct and Llama-3-70B-Instruct, respectively.
@@ C.2 Benchmarks
MathVista (lu2024mathvista) Testmini 1000 Accuracy
MathVision (mathvision) Full 3040 Accuracy
MathVerse (zhang2024mathverse) Vision-Only 788 Accuracy
DynaMath (zou2025dynamath) Full 5050 Accuracy
WeMath (wemath) Testmini 1740 Score (Strict)
LogicVista (xiao2024logicvista) Full 448 Accuracy
MMMU (yue2024mmmu) Validation 900 Accuracy
GSM8K (cobbe2021training) Test 1319 Accuracy
MATH (hendrycks_math) Test 5000 Accuracy
VisualProcessBench (wang2025visualprm) - 2866 F1-score
@@ C.3 Reward Choice
In this paper, we consider three reward choices of Athena-PRM for aggregating step scores into a final score (zhang2025lessons), including PRM-minium, PRM-last, and PRM-product. For all reasoning steps $\{a_{i}\}_{i=1}^{K}$, its corresponding reward for each step is $\{r_{i}\}_{i=1}^{K}$, the reward $r$ for each solution with different scoring methods are calculated as follows: 1) PRM-minium uses the minium reward across all steps, i.e., $r=\min\{r_{i}\}_{i=1}^{K}$; 2) PRM-last uses the reward of last step, i.e., $r=r_{K}$; 3) PRM-product uses the reward product of all steps, i.e. $r=\prod_{i=1}^{K}r_{i}$.
@@
Table 8: Best-of-8 performance comparison on the MathVista (lu2024mathvista) with Qwen2.5-VL-7B (qwen2.5vl) with different reward choices: last step reward, minium reward across all steps and product of all rewards. Default settings are marked in gray.
@@ Appendix D Case Study
Figure 4: A case study from VisualProcessBench (wang2025visualprm).
@@ 3. Angle properties:
Figure 5: A case study from VisualProcessBench (wang2025visualprm).
@@
Figure 6: A case study from MathVista (lu2024mathvista) with Qwen2.5-VL-72B (qwen2.5vl).
@@
Figure 7: A case study from MathVista (lu2024mathvista) with Qwen2.5-VL-72B (qwen2.5vl).